OpenStudy (anonymous):

Can someone please explain and show me? write the slope intercept equation for the line that passes through (-5,-8) and is Perpendicular to, 10x - 6y= -11

OpenStudy (anonymous):

\[10x - 6y= -11\] \[6y = 10x+11\] \[y = \frac53x+\frac{11}{6}\] Now perpendicular lines have inversive gradients. Thus the gradient of the new line is \[\frac35\] Thus our new line is: \[y = \frac35x+c\] Substitute in the new point \[-8 = \frac35(-5)+c\] \[-8 = -3+c\] \[c=-5\] And the equation is \[y = \frac35x+-5\]

OpenStudy (anonymous):

hi smart cow : )

OpenStudy (dumbcow):

agreed, except perpendicular lines must have opposite reciprocal slopes m = -3/5 hi loraaine

OpenStudy (dumbcow):

the slope of the given equation is 5/3 right? ........ the slope of perpendicular line must be opposite sign and reciprocal of that slope --> m = -3/5

OpenStudy (dumbcow):

follow rspencer's steps, except put the negative sign in there :)

OpenStudy (anonymous):

m=-3/5

OpenStudy (anonymous):

Oops, my bad

OpenStudy (dumbcow):

oh m is notation for slope of a line................"y=mx+b"

OpenStudy (anonymous):

thanks STAR~~MASTER!!